TROPYL RADICAL: A JAHN-TELLER SYSTEM WITH SEVEN-FOLD SYMMETRY

\underline{GYÖRGY TARCZAY},^{a} ILIAS SIOUTIS, TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The laser induced fluorescence (LIF) spectrum of the Ã^{2}E′′_{3} ← X̃^{2}E′′_{2} transition of the tropyl, C_{7}H_{7}, radical has been recorded and found to be in good agreement with the recently recorded REMPI spectrum.^{b} Dispersed fluorescence (DF) spectra, pumping several different transitions, have also been observed and yield the first experimental information on the ground-state vibronic levels. In order to analyze these spectra extensive ab inito calculations were carried out to determine the parameters of both the ground- and excited-state Jahn-Teller potential energy surfaces (PESs). Efforts to simulate both the LIF and DF spectra, and to refine the ab inito parameters by the aid of the SOCJT program will be discussed.^{c}

aPermanent address: Department of General and Inorganic Chemistry, Eötvös University, P.O. Box 32, H-1518 Budapest 112, Hungary

bT. Pino, F. Guthe, H. Ding, J. P. Maier J. Phys. Chem. A \underline{106}, 10022, 2002.

cT. A. Barkholtz, T. A. Miller Int. Rev. Phys. Chem. \underline{17}, 435, 1998.